OpenStudy (anonymous):

if F(x)=(x^2+1)/(3x-1), find F(-x)

OpenStudy (anonymous):

\[F(-x)=\frac{(-x)^2+1}{3(-x)-1}=\frac{x^2+1}{-3x-1}\]

OpenStudy (anonymous):

substitute (-x) everywhere you see an x, be sure to use parentheses:

OpenStudy (anonymous):

F(x)=(x^2+1)/(3x-1), find F(-x) F(-x)=((-x)^2+1)/(3(-x)-1) =(x^2+1)/(-3x-1)

OpenStudy (anonymous):

Thanks.